OpenStudy (anonymous):

Given f(x) = 8x + 13 and g(x) = x^2 - 5x, find the function (f + g)(x). Use the ^ key for the exponent. Enter the answer as the following example: -5x^2+4x-8

OpenStudy (anonymous):

similar to the previous one, but this time there is no annoying minus sign so you can compute \[f(x)+g(x) = 8x+13+x^2-5x\] in one step by combining like terms

OpenStudy (anonymous):

the only like terms are \(8x\) and \(-5x\) so your answer should be \[(f+g)(x)=x^2+3x+13\]

OpenStudy (anonymous):

ohh ok, i get it now, lol thanks.

OpenStudy (anonymous):

now i have a question is there really a "cactus high"?

OpenStudy (anonymous):

lol ya

OpenStudy (anonymous):

wow!

OpenStudy (anonymous):

lol